# MetaML-Pro: Cross-Stage Design Flow Automation for Efficient Deep Learning Acceleration

### v2

@@ Abstract.
This paper presents a unified framework for codifying and automating optimization strategies to efficiently deploy deep neural networks (DNNs) on resource-constrained hardware, such as FPGAs, while maintaining high performance, accuracy, and resource efficiency. Deploying DNNs on such platforms involves addressing the significant challenge of balancing performance, resource usage (e.g., DSPs and LUTs), and inference accuracy, which often requires extensive manual effort and domain expertise. Our novel approach addresses two core key issues: (i) encoding custom optimization strategies and (ii) enabling cross-stage optimization search. In particular, our proposed framework seamlessly integrates programmatic DNN optimization techniques with high-level synthesis (HLS)-based metaprogramming, leveraging advanced design space exploration (DSE) strategies like Bayesian optimization to automate both top-down and bottom-up design flows. Hence, we reduce the need for manual intervention and domain expertise. In addition, the framework introduces customizable optimization, transformation, and control blocks to enhance DNN accelerator performance and resource efficiency. Experimental results demonstrate up to a 92% DSP and 89% LUT usage reduction for select networks, while preserving accuracy, along with a 15.6-fold reduction in optimization time compared to grid search. These results highlight the potential for automating the generation of resource-efficient DNN accelerator designs with minimum effort.
FPGAs, algorithm-hardware codesign, design automation, design space exploration
@@ 1. Introduction
The field of deep learning has witnessed unprecedented growth in recent years, driven by the increasing demand for efficient and high-performance applications (LeCun et al., 2015). Consequently, FPGA-based deep neural network (DNN) accelerator design and optimization have gained significant attention (Zhang et al., 2015, 2020). The development of an efficient FPGA-based DNN design requires a diverse skill set that combines expertise in machine learning with low-level knowledge of the target hardware architecture (Sze et al., 2017). Optimizing these DNN designs is a complex process, as it involves balancing competing objectives. On one hand, high accuracy during inference is crucial from an application perspective. On the other hand, the design must be optimized for the underlying hardware architecture, meeting power, latency and throughput requirements while fitting into FPGA devices (Duarte et al., 2018; Coelho et al., 2021). Achieving an optimal balance between these conflicting objectives requires careful consideration and effective optimization strategies (Wang et al., 2019).
Existing optimization techniques for DNNs and hardware have a limitation in that they are often handled in separate stages, with information exchanged only in a top-down manner from the application to hardware. Bottom-up information flow, where hardware design insights inform DNN model optimization, typically requires manual intervention. However, this information exchange across multiple stages can significantly enhance the optimization process, as application and hardware optimizations can interact and affect each other. For instance, modifying the DNN model architecture could impact the FPGA resource utilization, and optimizing the FPGA design could influence the DNN model’s accuracy (Wang et al., 2019). In addition, there has been limited focus on combining optimization strategies targeting different abstraction levels, such as algorithmic level neural networks and High Level Synthesis (HLS) C++, hindering the selection of the most effective combination of optimization techniques for a given problem. In summary, few studies haven thoroughly explored the vast configuration space that emerges when combining DNN optimization techniques across both software and hardware domains.
@@
Our proposed co-optimization framework, as illustrated in Fig. 1, operates across various computational spaces, such as software and High-Level Synthesis (HLS), and can potentially target diverse hardware platforms including FPGAs and ASICs (see C1). Each computational space represents an autonomous, custom optimisation task (e.g., OPT-1 to OPT-5, and potentially more), governed by tunable parameters. Moreover, this work utilizes a Bayesian (Reagen et al., 2017; Mehrabi et al., 2020; Tuli et al., 2023) guided nested-loop optimization structure involving both local and global phases, combining very distinct optimization tasks and incorporating feedback from the latter stages of the design flow to refine the process iteratively (see C2).
@@
  • A novel co-optimization framework that enables modular and adaptive hardware-software exploration for FPGA-based DNN accelerators. It independently encodes and combines optimizations for efficient design refinement. A cross-stage optimization search automates selection, ordering, and tuning across abstraction levels, dynamically adapting to hardware and application constraints through top-down and bottom-up strategies, ensuring efficient and high-performance designs (Section 3).
  • A library of reusable optimization, transformation, and control tasks designed to be customizable and flexible, and that can be easily integrated into our co-optimization framework. This library includes optimization tasks such as pruning, quantization, and scaling to enhance DNN accelerator performance and resource efficiency. Moreover, we leverage metaprogramming to programmatically analyze and modify HLS C++ designs, enhancing design flexibility, performance tuning, and resource efficiency. In addition, we integrate Bayesian optimization to accelerate design space exploration while maintaining an optimal balance between accuracy and hardware efficiency. Some of the tasks in our library are specific to certain applications and target technologies, while others remain agnostic, providing versatility and adaptability within the framework (Section 4).
@@
This paper expands on our previous studies (Que et al., 2023, 2024a, 2024c) which primarily focus on local optimizations. In (Que et al., 2023), we propose MetaML, an optimization framework, that customizes design flow for DNNs with local optimizations. In (Que et al., 2024a), we introduce control capabilities within the optimization framework, while (Que et al., 2023) focuses on Quantization Heuristic Search (QHS) optimization task for compressing DNN accelerators. However, these prior studies are limited in their ability to perform global optimization due to their focus on isolated optimization tasks. The inability to account for interdependencies between optimization tasks across stages mean that the overall design might not have achieved the optimal balance between accuracy, performance, and hardware resource. This work addresses these limitations by incorporating bottom-up feedback and leveraging Bayesian optimization to enable global optimization across the entire design flow. By adjusting optimization tasks at each stage based on performance feedback, our framework ensures that each optimization choice contributes to a more cohesive and globally optimal solution. This extended approach maintains high accuracy while improving resource efficiency, specifically in resource-constrained environments. We plan to open-source MetaML-Pro to the research community, aiming to inspire further research and development in this area.
@@ 2. Related Work
Other approaches offer end-to-end software frameworks, such as Xilinx’s Vitis AI (Kathail, 2020) and Intel’s OpenVINO (Ope, 2023), which optimize DNNs with pre-built optimizations for deployment on specific target technologies. However, they are not designed for easy addition of new optimization strategies or to search for the most effective combined strategies.
Furthermore, some frameworks allow developers to describe and customize DNN optimization strategies, but their scope is limited. For instance, ScaleHLS (Ye et al., 2022) and SOTA-OPT (Agostini et al., 2022) focus on HLS-based optimization strategies and hardware designs, but their optimization scope is restricted to MLIR. TVM (Chen et al., 2018) is a general-purpose DNN compilation framework that offers performance portability across different types of devices, but optimizations occur solely at the graph (IR) level.
Moreover, frameworks, such as FINN (Umuroglu et al., 2017), HLS4ML (Duarte et al., 2018), and fpgaConvNet (Venieris and Bouganis, 2016), provide optimized hardware building blocks for FPGA-based DNN accelerators, and allow optimization strategies to be codified using these blocks. However, they do not support automated bottom-up optimization flows, where hardware stage information guides the optimization of the application (DNN) stage.
@@ 3.1. Overview
Figure 2. A typical FPGA design flow. This FPGA design flow begins with defining the specifications (SPEC), followed by implementing and testing the design in software (SW). If High-Level Synthesis (HLS) is used, high-level code (e.g., C/C++) is converted into hardware description language (HDL). The design then moves to the Register Transfer Level (RTL) stage, where it is described in HDL (e.g., Verilog/VHDL) and synthesized into a netlist of logic gates. Finally, the netlist is used to generate a bitstream, which is loaded onto the FPGA to configure the hardware. Each stage progressively refines the design from concept to implementation.
@@ 3.4. Building a Strategy
Listing.1 provides the Python code for implementing the pruning strategy depicted in Fig.8(a) and discussed in Section 5.2. The code consists of three sections: (i) lines 4–9 builds the design-flow architecture graph by connecting the corresponding tasks. The >>much-greater-than$>$ > and <<much-less-than$<$ < operators are utilized to define the connections between task instances, and the default name of each task instance serves as its identifier, although it can be overridden by providing it as an argument (line 6). The end result is a cyclic-directed graph. (ii) lines 11–23 define the configuration object, which holds all the parameters for configuring the tasks. Parameters referred to as TaskType::parameter are passed to all tasks of that task type (lines 16–19), whereas parameter names in the format of InstanceName@parameter are passed to a specific instance (line 15). All other parameters are global (lines 20–21) and accessible to any task. Finally, (iii) line 25 executes the design-flow by passing the configuration object.
The execution of the design-flow starts with graph validation, which ensures that there is at least one source task and no multiplicity constraints are violated. Tasks are executed by an internal scheduler using a thread pool, which assigns jobs to available workers or stalls until a worker becomes available. After validating the architecture graph, an empty meta-model is generated using the supplied configuration, and the scheduler runs the source task. When a task completes execution, it submits a job to execute the next set of tasks in the flow. The STOP task terminates execution when all submitted jobs are completed. Modifying the design-flow architecture and updating its configuration parameters can lead to new optimization strategies, which is discussed next.
@@ 4.2. Auto-Pruning with Binary Search
Starting at 0% pruning rate, the auto-pruning algorithm obtains initial accuracy $Acc_{p0}$ at step 1 (s1). It then uses a binary search approach, increasing or decreasing the pruning rate based on whether the accuracy loss is within a user-defined tolerance ($\leq\alpha_{p}$) as shown in Table 1. The algorithm terminates when the rate difference is below a threshold ($\beta_{p}$). The number of steps is determined by $1+log_{2}(1/\beta_{p})$. Algorithm search steps and direction are shown in Fig. 9. Both the tolerance ($\leq\alpha_{p}$) and threshold ($\beta_{p}$) values are initially set to 2% in this work.
@@ 4.4. Auto-Quantization with Heuristic Search
In our framework, we introduce quantization task based on the Quantization Heuristic Search (QHS) algorithm (Que et al., 2024c). It automatically determines the lowest bitwidth for each layer while maintaining a given level of accuracy loss. Formally, the objective of this task is defined as:
@@
Fig. 5 shows the steps of the proposed QHS algorithm. The optimization task begins by accepting a model with a maximum tolerable accuracy loss, denoted as $\alpha_{q}$. Initially, virtual layers are created, and lossless reduction is performed by minimizing the integer bits representing model parameters using the base-2 logarithm of the largest weight and bias in each layer, plus one bit for the sign to avoid saturation. Subsequently, all parameters (weights, biases and results) are initially assumed to be further reducible. The total bit-widths of these parameters are then reduced by one bit if they are considered reducible. This quantization is applied to the C++ kernel, and the accuracy is evaluated through a runtime simulation. If the accuracy loss remains within the user-specified tolerance, $\alpha_{q}$, this step is repeated. If the accuracy loss exceeds the tolerance, the algorithm identifies and blocks non-reducible precisions. It determines which precisions are most sensitive by testing the reduction impact on accuracy. If a bit-width reduction breaks the accuracy constraint, that precision is marked non-reducible. The process repeats until no further reductions are possible within the accuracy constraints.
@@ 4.5. Metaprogramming for HLS Design Optimization
A novel aspect of our co-optimization approach is the integration of metaprogramming into hardware design optimization. In this section, we describe its key motivation and benefits, and how we integrate metaprogramming in our co-optimization framework.
Limitations of template-based approaches. Traditional FPGA-based DNN acceleration frameworks typically rely on manually crafted templates to generate hardware designs. These templates encode expert knowledge, ensuring optimized implementations for a given set of parameters as illustrated in Fig. 6(a). However, they impose inherent limitations in flexibility, adaptability, and scalability. First, templates generate only a limited number of design variants by adjusting predefined parameters, restricting design exploration. They are optimized for specific architectures and are not easily generalizable to new DNN models or hardware platforms.
Extending these templates to support new DNN architectures requires deep domain knowledge in hardware design, making the process resource-intensive. Hardware engineers must manually introduce modifications, which increases development time and effort. Additionally, the modifications are constrained to specific parameterized components, preventing broader architectural transformations that might be necessary for emerging hardware and application demands. As a result, template-based methods struggle to efficiently explore the vast design space available in FPGA-based DNN acceleration.
Benefits of metaprogramming. Our metaprogramming approach addresses these challenges by dynamically analyzing and modifying HLS C++ source code using Python. This technique allows greater flexibility in exploring hardware optimization strategies. One key advantage is the ability to automate the insertion and tuning of HLS pragmas, such as loop unrolling and pipelining, without requiring manual intervention. By adjusting these pragmas dynamically, the hardware implementation can be fine-tuned for improved performance and resource utilization.
Additionally, metaprogramming facilitates code restructuring to improve computational efficiency and reduce design complexity. By programmatically modifying code, redundant computations can be eliminated, and data dependencies can be optimized. Another important capability is automated data type adjustments for quantization, which optimize memory footprint and computational precision by dynamically adapting numerical representations. Unlike static templates, which rely on predefined precision settings, metaprogramming can analyze the computational requirements of different layers and assign appropriate precision dynamically.
Another significant benefit is the incorporation of specialized FPGA libraries (Vandebon et al., 2021), enhancing performance for target platforms. By integrating device-specific optimizations at the source-code level, our approach ensures that the generated hardware designs are optimized for the underlying FPGA architecture. This adaptability makes metaprogramming a powerful tool for hardware-aware DNN optimization.
@@
Example of an HLS metaprogram. Fig. 6(b) illustrates an HLS metaprogram that dynamically modifies a given C++ source file. The process begins with the construction of an Abstract Syntax Tree (AST) representation of the C++ code. The AST provides a structured representation of the source code, allowing precise modifications for any of its constructs. The metaprogram then performs code querying and analysis to identify outermost loop structures within kernel functions. These loops are critical for performance optimization, as their execution directly impacts resource utilization and latency.

### v3

@@ Abstract.
This paper presents a unified framework for codifying and automating optimization strategies to efficiently deploy deep neural networks (DNNs) on resource-constrained hardware, such as FPGAs, while maintaining high performance, accuracy, and resource efficiency. Deploying DNNs on such platforms involves addressing the significant challenge of balancing performance, resource usage (e.g., DSPs and LUTs), and inference accuracy, which often requires extensive manual effort and domain expertise. Our novel approach addresses two core issues: (i) encoding custom optimization strategies and (ii) enabling cross-stage optimization search. In particular, our proposed framework seamlessly integrates programmatic DNN optimization techniques with high-level synthesis (HLS)-based metaprogramming, leveraging advanced design space exploration (DSE) strategies like Bayesian optimization to automate both top-down and bottom-up design flows. Hence, we reduce the need for manual intervention and domain expertise. In addition, the framework introduces customizable optimization, transformation, and control blocks to enhance DNN accelerator performance and resource efficiency. We further formalize a cross-stage, constrained Bayesian optimization procedure that couples predicate–action bottom-up feedback (via BRANCH) with FORK/REDUCE order search, enabling automated selection, ordering, and tuning across software and HLS tasks. Experimental results demonstrate up to a 92% DSP and 89% LUT usage reduction for select networks, while preserving accuracy, along with a 15.6-fold reduction in optimization time compared to grid search. These results highlight the potential for automating the generation of resource-efficient DNN accelerator designs with minimal effort, resulting in large resource savings with bounded exploration cost.
FPGAs, algorithm-hardware codesign, design automation, design space exploration, MetaML
@@ 1. Introduction
The field of deep learning has witnessed unprecedented growth in recent years, driven by the increasing demand for efficient and high-performance applications (LeCun et al., 2015; Que et al., 2021b; Wojcicki et al., 2022; Lou et al., 2025). Consequently, FPGA-based deep neural network (DNN) accelerator design and optimization have gained significant attention (Que et al., 2022b, 2020c, 2020b, c). The development of an efficient FPGA-based DNN design requires a diverse skill set that combines expertise in machine learning with low-level knowledge of the target hardware architecture (Sze et al., 2017; Que et al., 2020a). Optimizing these DNN designs is a complex process, as it involves balancing competing objectives (Que et al., 2025a). On one hand, high accuracy during inference is crucial from an application perspective. On the other hand, the design must be optimized for the underlying hardware architecture, meeting power, latency, and throughput requirements while fitting into FPGA devices (Que et al., 2024b, 2025b). Achieving an optimal balance between these conflicting objectives requires careful consideration and effective optimization strategies (Que et al., 2025a).
Existing optimization techniques for DNNs and hardware have a limitation in that they are often handled in separate stages, with information exchanged only in a top-down manner from the application to hardware. Bottom-up information flow, where hardware design insights inform DNN model optimization, typically requires manual intervention. However, this information exchange across multiple stages can significantly enhance the optimization process, as application and hardware optimizations can interact and affect each other. For instance, modifying the DNN model architecture could impact the FPGA resource utilization, and optimizing the FPGA design could influence the DNN model’s accuracy (Wang et al., 2019; Que et al., 2024b). In addition, there has been limited focus on combining optimization strategies targeting different abstraction levels, such as algorithmic level neural networks and High Level Synthesis (HLS) C++, hindering the selection of the most effective combination of optimization techniques for a given problem. In summary, only a limited number of studies have thoroughly explored the vast configuration space that emerges when combining DNN optimization techniques across both software and hardware domains.
@@
Our proposed co-optimization framework, as illustrated in Fig. 1, operates across various computational spaces, such as software and High-Level Synthesis (HLS), and can target diverse FPGA platforms (see C1). Each computational space represents an autonomous, custom optimisation task (e.g., OPT-1 to OPT-5, and potentially more), governed by tunable parameters. Moreover, this work utilizes a Bayesian (Mehrabi et al., 2020; Tuli et al., 2023; Que et al., 2025a) guided nested-loop optimization structure involving both local and global phases, combining very distinct optimization tasks and incorporating feedback from the latter stages of the design flow to refine the process iteratively (see C2). Finally, while we instantiate Bayesian optimization for the outer DSE, the controller is algorithm-agnostic and can host evolutionary or RL-based search as drop-in alternatives.
@@
  • A novel co-optimization framework that enables modular and adaptive hardware-software exploration for FPGA-based DNN accelerators. It independently encodes and combines optimizations for efficient design refinement.
  • A cross-stage optimization algorithm. We formalize a nested-loop, constraint-aware Bayesian optimization procedure over software- and HLS-level optimization tasks. The algorithm uses predicate–action bottom-up feedback (via BRANCH) to adapt accuracy/utilization tolerances on constraint violations, and FORK/REDUCE to search task permutations and aggregate by Pareto dominance. This automates selection, ordering, and tuning across abstraction levels.
  • A library of reusable optimization, transformation, and control tasks designed to be customizable and flexible, and that can be easily integrated into our co-optimization framework. This library includes optimization tasks such as pruning, quantization, and scaling to enhance DNN accelerator performance and resource efficiency. Moreover, we leverage metaprogramming to programmatically analyze and modify HLS C++ designs, enhancing design flexibility, performance tuning, and resource efficiency. Some of the tasks in our library are specific to certain applications and target technologies, while others remain agnostic, providing versatility and adaptability within the framework.
@@
This paper expands on our previous studies (Que et al., 2023, 2024a, 2024d) which primarily focus on local optimizations. In (Que et al., 2023), we propose MetaML, an optimization framework, that customizes design flow for DNNs with local optimizations. In (Que et al., 2024a), we introduce control capabilities within the optimization framework, while (Que et al., 2023) focuses on Quantization Heuristic Search (QHS) optimization task for compressing DNN accelerators. However, these prior studies are limited in their ability to perform global optimization due to their focus on isolated optimization tasks. The inability to account for interdependencies between optimization tasks across stages mean that the overall design might not have achieved the optimal balance between accuracy, performance, and hardware resource. This work addresses these limitations by incorporating bottom-up feedback and leveraging Bayesian optimization to enable global optimization across the entire design flow. By adjusting optimization tasks at each stage based on performance feedback, our framework ensures that each optimization choice contributes to a more cohesive and globally optimal solution. This extended approach maintains high accuracy while improving resource efficiency, specifically in resource-constrained environments.
@@ 2. Related Work
Other approaches offer end-to-end software frameworks, such as Xilinx’s Vitis AI (Kathail, 2020) and Intel’s OpenVINO (31), which optimize DNNs with pre-built optimizations for deployment on specific target technologies. However, they are not designed for easy addition of new optimization strategies or to search for the most effective combined strategies.
Furthermore, some frameworks allow developers to describe and customize DNN optimization strategies, but their scope is limited. For instance, ScaleHLS (Ye et al., 2022) and SOTA-OPT (Agostini et al., 2022) focus on HLS-based optimization strategies and hardware designs, but their optimization scope is restricted to MLIR. TVM (Chen et al., 2018) is a general-purpose DNN compilation framework that offers performance portability across different types of devices, but optimizations occur solely at the graph (IR) level. da4ml (Sun et al., 2025b) is an open-source library that compiles machine-learning models into hardware designs. It targets fully pipelined implementations with an initiation interval (II) of 1, applies distributed arithmetic (DA) optimizations, and can emit synthesizable Verilog and VHDL for multiple FPGA back-ends.
Moreover, frameworks, such as FINN (Umuroglu et al., 2017), HLS4ML (Schulte et al., 2025), and fpgaConvNet (Venieris and Bouganis, 2016), provide optimized hardware building blocks for FPGA-based DNN accelerators, and allow optimization strategies to be codified using these blocks. However, they do not support automated bottom-up optimization flows, where hardware stage information guides the optimization of the application (DNN) stage.
@@ 3.1. Overview
Figure 2. A typical FPGA design flow. This FPGA design flow begins by defining the specifications (SPEC), followed by implementing and testing the design in software (SW). If High-Level Synthesis (HLS) is used, high-level code (e.g., C/C++) is converted into hardware description language (HDL). The design then moves to the Register Transfer Level (RTL) stage, where it is described in HDL (e.g., Verilog/VHDL) and synthesized into a netlist of logic gates. Finally, the netlist is used to generate a bitstream, which is loaded onto the FPGA to configure the hardware. Each stage progressively refines the design from concept to implementation.
@@ 3.4. Building a Strategy
Listing 1 provides the Python code for implementing the pruning strategy depicted in Fig.9(a) and discussed in Section 5.2. The code consists of three sections: (i) lines 4–9 builds the design-flow architecture graph by connecting the corresponding tasks. The >>>> and <<<< operators are utilized to define the connections between task instances, and the default name of each task instance serves as its identifier, although it can be overridden by providing it as an argument (line 6). The end result is a cyclic-directed graph. (ii) lines 11–23 define the configuration object, which holds all the parameters for configuring the tasks. Parameters referred to as TaskType::parameter are passed to all tasks of that task type (lines 16–19), whereas parameter names in the format of InstanceName@parameter are passed to a specific instance (line 15). All other parameters are global (lines 20–21) and accessible to any task. Finally, (iii) line 25 executes the design-flow by passing the configuration object.
The execution of the design-flow starts with graph validation, which ensures that there is at least one source task and no multiplicity constraints are violated. Tasks are executed by an internal scheduler using a thread pool, which assigns jobs to available workers or stalls until a worker becomes available. After validating the architecture graph, an empty meta-model is generated using the supplied configuration, and the scheduler runs the source task. When a task completes execution, it submits a job to execute the next set of tasks in the flow. The STOP task terminates execution after all submitted jobs are completed. Modifying the design-flow architecture and updating its configuration parameters can lead to new optimization strategies, which is discussed next.
[…]
@@ 4.2. Auto-Pruning with Binary Search
Starting at 0% pruning rate, the auto-pruning algorithm obtains initial accuracy $Acc_{p0}$ at step 1 (s1). It then uses a binary search approach, increasing or decreasing the pruning rate based on whether the accuracy loss is within a user-defined tolerance ($\leq\alpha_{p}$) as shown in Table 1. The algorithm terminates when the rate difference is below a threshold ($\beta_{p}$). The number of steps is determined by $1+log_{2}(1/\beta_{p})$.
@@ 4.4. Auto-Quantization with Heuristic Search
In our framework, we introduce quantization task based on the Quantization Heuristic Search (QHS) algorithm (Que et al., 2024d), a post-quantization method. It automatically determines the lowest bitwidth for each layer while maintaining a given level of accuracy loss. Formally, the objective of this task is defined as:
@@
Fig. 5 shows the steps of the proposed QHS algorithm. The optimization task begins by accepting a model with a maximum tolerable accuracy loss, denoted as $\alpha_{q}$. Initially, virtual layers are created, and lossless reduction is performed by minimizing the integer bits representing model parameters using the base-2 logarithm of the largest weight and bias in each layer, plus one bit for the sign to avoid saturation. Subsequently, all parameters (weights, biases and results) are initially assumed to be further reducible. The total bit-widths of these parameters are then reduced by one bit if they are considered reducible. This quantization is applied to the C++ kernel, and the accuracy is evaluated through a runtime simulation. If the accuracy loss remains within the user-specified tolerance, $\alpha_{q}$, this step is repeated. If the accuracy loss exceeds the tolerance, the algorithm identifies and blocks non-reducible precisions. It determines which precisions are most sensitive by testing the reduction impact on accuracy. If a bit-width reduction breaks the accuracy constraint, that precision is marked non-reducible. The process repeats until no further reductions are possible within the accuracy constraints. Support for Quantization-aware training, such as HGQ (Sun et al., 2026), is left to future work.
@@ 4.5. Metaprogramming for HLS Design Optimization
A novel aspect of our co-optimization approach is the integration of metaprogramming into hardware design optimization, which enables the implementation of the Quantization Heuristic Search (QHS) algorithm described in Section 4.4. In this section, we describe the key motivation and benefits of metaprogramming, and how we integrate it in our co-optimization framework.
Beyond template-based and code generator approaches. Template-based frameworks and code generator modifications (for instance, extending HLS4ML) are effective when optimization choices can be captured as a fixed set of global parameters. For example, templates can expose settings such as default precision, loop unroll factors, or hardware interface options. Similarly, modifying a code generator allows these parameters to be realized directly in generated code. However, both approaches share two fundamental limitations:
  • Restricted scope of changes: Templates and generators can only express transformations that were foreseen and encoded by their designers. Adding new optimisation dimensions requires extending the generator itself, which is costly and often tool-specific.
  • Limited applicability across sources: In practice, real workflows mix generated code (e.g., HLS4ML) with user-written kernels, vendor libraries, and legacy C++. Template extensions or generator patches cannot easily propagate optimisations across such heterogeneous sources.
Benefits of metaprogramming. Our metaprogramming approach addresses these limitations by operating directly on the C++ Abstract Syntax Tree (AST). This enables:
  • Local, structural optimizations such as loop-specific pragma placement, which templates cannot expose without proliferating knobs.
  • Cross-kernel consistency edits, for example rewriting all types in a conv → pool → batch norm cascade to ensure coherent quantisation without re-generation.
  • Device-aware refactorings guided by design-space exploration, where metaprograms inject vendor-specific constructs based on feedback.
Quantization illustrates this distinction clearly. A template or generator can reduce precision globally (e.g., set all layers to 8-bit), but cannot ensure type consistency across intermediate streams or flexibly adapt precision to device constraints without modifying the generator itself. Our metaprogram, by contrast, queries the AST to identify cascaded kernels, rewrites typedefs and template arguments in one pass, and produces a coherent low-precision dataflow candidate automatically.
In summary, templates and generators are well-suited for global configuration, while metaprogramming is essential when optimisation requires local, structural, or device-specific edits. This generality allows the same refactoring (e.g., quantisation) to be applied not only to HLS4ML output, but equally to handwritten or third-party C++ code, demonstrating the broader utility of our approach.
@@
Example of an HLS metaprogram. Fig. 6(b) illustrates an HLS metaprogram that dynamically modifies a given C++ source file using Python. The process begins with the construction of an Abstract Syntax Tree (AST) representation of the C++ code. The AST provides a structured representation of the source code, allowing precise modifications for any of its constructs. The metaprogram then performs code querying and analysis to identify outermost loop structures within kernel functions. These loops are critical for performance optimization, as their execution directly impacts resource utilization and latency.